# Constrained mean-field control with singular control: Existence, stochastic maximum principle and constrained FBSDE

### v2

@@ 5 Uniqueness and Stability of Solution to Constrained FBSDE
Hence, combining the above two inequalities leads to $\bar{X}=0,\bar{Y}=0$ and $\bar{Z}=0$. It then follows that $\bar{\alpha}=0$ and $\int_{0}^{\cdot}G(s)\mathrm{d}\bar{\zeta}_{s}=0$. Moreover, we have $\int_{0}^{t}A(s)\bar{\eta}_{s}\mathrm{d}s+\int_{0}^{t}B(s)\mathbb{E}\left[\bar{\eta}_{s}\right]\mathrm{d}s=0$ for all $t\in[0,T]$. Taking expectations on both sides of the above result, together with (C1) in Assumption 3, we conclude that $\mathbb{E}\left[\bar{\eta}_{t}\right]=0$, for $m$-a.e. $t\in[0,T]$, and hence $\int_{0}^{t}A(s)\bar{\eta}_{s}\mathrm{d}s=0$, for $m$-a.e. $t\in[0,T]$. This indicates that $\bar{\eta}=0$ in $\mathbb{L}_{\mathbb{F}}^{2}$ by utilizing Assumption 3-(C1). ∎
Remark 5.5.
Here, we only consider the case where $Y_{T}=g(X_{T},\mu_{T})$ with $g=0$ in the constrained FBSDE (5)-(27). However, the uniqueness result still holds if we additionally assume:
  • (C5)
    Set $\Psi(x,\mu):=\nabla_{x}g(x,\mu)+\int_{\mathbb{R}^{n}}\partial_{\mu}g(x^{\prime},\mu)(x)\mu(\mathrm{d}x^{\prime})$ for $(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$. Then, $\Psi$ satisfies the monotone condition $(\Psi(x_{1},\mu)-\Psi(x_{2},\mu))(x_{1}-x_{2})\geq 0$ and $\int_{\mathbb{R}^{n}}(\Psi(x,\mu_{1})-\Psi(x,\mu_{2}))x^{\prime}(\mu_{1}-\mu_{2})(\mathrm{d}x^{\prime})\geq 0$ for $x,x_{1},x_{2}\in\mathbb{R}^{n}$ and $\mu,\mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R}^{n})$.
@@
    The constrained function $\phi$ is independent of the control variable and admits the linear form that $\phi(t,x,\mu)=A(t)x+B(t)\int_{\mathbb{R}^{n}}x\mu(\mathrm{d}x)+C(t)$ for $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$. Here, the coefficients $A,B,C\in C([0,T],\mathbb{R}^{n})$ satisfying $|A(t)|\geq c_{0}>0$ and $|A(t)+B(t)|\geq c_{0}>0$ for some $c_{0}>0$ independent of $t\in[0,T]$.
@@
Let Assumption 2, (C1’), (C3)-(C4) of Assumption 3 hold and $\xi_{1},\xi_{2}\in L^{2}((\Omega,\mathcal{F}_{0},\mathbb{P});\mathbb{R}^{n})$. Denote by $(\widehat{X}^{i},\widehat{\alpha}_{t}^{i},\widehat{\zeta}^{i}_{t},Y^{i},Z^{i},\eta^{i})$ the unique solutions to the FBSDE (5)-(27) with respective initial data $\xi_{i}$. Set $\bar{X}_{t}:=X_{t}^{1}-X_{t}^{2}$, $\bar{\zeta}_{t}:=\zeta_{t}^{1}-\zeta_{t}^{2}$, $\bar{Y}_{t}:=Y_{t}^{1}-Y_{t}^{2}$, $\bar{Z}_{t}:=Z_{t}^{1}-Z_{t}^{2}$ and $\bar{\eta}_{t}:=\eta_{t}^{1}-\eta_{t}^{2}$ for $t\in[0,T]$. Then, there exists a constant $C>0$ depending on $(T,M,\alpha,\beta)$ only such that
@@
$\qquad\qquad\leq\mathbb{E}\left[|\xi_{1}-\xi_{2}|^{2}\right].$ (31)
@@
Recall $\Delta b(t),\Delta\sigma(t),\Delta F(t)$ defined in the proof of Theorem 5.4. Then, one has
$\int_{0}^{t}G(s)\mathrm{d}\bar{\zeta}_{s}=\bar{X}_{t}-(\xi_{1}-\xi_{2})-\int_{0}^{t}\Delta b(s)\mathrm{d}s-\int_{0}^{t}\Delta\sigma(s)\mathrm{d}s.$ (32)
@@
$\left|\bar{Y}_{t}\right|^{2}+\int_{t}^{T}|\bar{Z}_{s}|^{2}\mathrm{d}s$ $=2\int_{t}^{T}\bar{Y}_{s}^{\top}\Delta F(s)\mathrm{d}s-2\int_{t}^{T}\bar{Y}_{s}\bar{Z}_{s}\mathrm{d}W_{s}-2\int_{t}^{T}\bar{\eta}_{s}A(s)^{\top}\bar{Y}_{s}\mathrm{d}s$
@@
Using Assumption 3-(C4), the condition (C1’), the BDG’s inequality and the Gronwall’s inequality, we can similarly derive, for some constant $C>0$ depending on $(M,\epsilon,T)$ only,
@@
Applying integration by parts to $\bar{X}_{t}^{\top}\bar{Y}_{t}$ from $t=0$ to $T$, we have
@@
Taking expectations on both sides of the above equality and following the proof of Theorem 5.4, we deduce that
$-\mathbb{E}\left[\bar{Y}_{0}^{\top}\left(\xi^{1}-\xi^{2}\right)\right]+\beta\mathbb{E}\left[\int_{0}^{T}(|\bar{X}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2})\mathrm{d}t\right]\leq 0.$ (37)
As a result, the inequality (37) leads to that
$\alpha\mathbb{E}\left[\left|\bar{X}_{T}\right|^{2}\right]+\beta\mathbb{E}\left[\int_{0}^{T}(|\bar{X}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2})\mathrm{d}t\right]$ $\leq\mathbb{E}\left[\bar{Y}_{0}^{\top}(\xi^{1}-\xi^{2})\right]\leq\frac{2}{\epsilon}\mathbb{E}\left[\left|\bar{Y}_{0}\right|^{2}\right]$
$\quad+\frac{1}{2\epsilon}\mathbb{E}\left[\left|\xi^{1}-\xi^{2}\right|^{2}\right].$
@@
$\mathbb{E}\left[\left|\bar{X}_{T}\right|^{2}+\int_{0}^{T}\left(|\bar{X}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2}\right)\mathrm{d}t\right]\leq C\mathbb{E}\left[\left|\xi^{1}-\xi^{2}\right|^{2}\right].$ (38)
Combining the Lipschitz continuity of $\hat{u}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto U\subset\mathbb{R}^{l}$ in Assumption 3-(C3), (32), (35) and the condition (C1’), we conclude the stability (5.6) from (38) above. ∎
@@
When considering $Y_{T}=g(X_{T},\mu_{T})$, we further assume that
  • (C5’)
    Set $\Psi(x,\mu):=\nabla_{x}g(x,\mu)+\int_{\mathbb{R}^{n}}\partial_{\mu}g(y,\mu)(x)\mu(\mathrm{d}y)$ for $(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$. The function $\Psi$ is Lipschitz continuous and satisfies the coercivity condition that, for some $\alpha,M>0$ and $(x^{1},\mu^{1},X^{1}),(x^{2},\mu^{2},X^{2})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times L^{2}((\Omega,\mathcal{F}_{T},\mathbb{P}),\mathbb{R}^{n})$,
    $|\Psi(x^{1},\mu^{1})-\Psi(x^{2},\mu^{2})|\leq M(|x^{1}-x^{2}|+\mathcal{W}_{2,\mathbb{R}^{n}}(\mu^{1},\mu^{2})),$
In addition to the setting in Theorem 5.6, let $Y^{i}$ have terminal value $g^{i}(X_{T}^{i},\mu_{T}^{i})$ and $g^{1},g^{2}\in C^{1;1}(\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n});\mathbb{R})$ with $g^{1}$ satisfying the condition (C5’). Then, it holds that
@@
$\qquad\qquad\qquad\leq C\left\{\mathbb{E}\left[|\xi_{1}-\xi_{2}|^{2}\right]+\sup_{(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})}\left|\Psi^{1}(x,\mu)-\Psi^{2}(x,\mu)\right|^{2}\right\}$
@@
Acknowledgements L. Bo and J. Wang are supported by National Natural Science of Foundation of China (No. 12471451), Natural Science Basic Research Program of Shaanxi (No. 2023-JC-JQ-05), Shaanxi Fundamental Science Research Project for Mathematics and Physics (No. 23JSZ010). X. Yu and J. Wang are supported by the Hong Kong RGC General Research Fund (GRF) under grant No. 15214125 and by the Hong Kong Polytechnic University research grant under No. P0045654.
@@
This, together with Lemma 3.2, indicates that $\mathbb{E}^{R}[\int_{0}^{T}|\alpha_{t}|^{p}\mathrm{d}t]<\infty$. On the other hand, using the construction of $\alpha$ again, we deduce that, for any $g\in C_{b}^{2}(\mathbb{R}^{n})$ and $(t,\omega)\in[0,T]\times\Omega$,
$g(X_{t}(\omega))-\int_{0}^{t}\mathbb{G}g(s,X_{s}(\omega),\mathcal{L}^{R}(X_{s}),\alpha_{s}(\omega))\mathrm{d}s-\int_{0}^{t}\nabla_{x}g(X_{s}(\omega))G(s)\mathrm{d}\zeta_{s}(\omega)$
$\qquad-\sum_{0<s\leq t}\left(g(X_{s}(\omega))-g(X_{s-}(\omega))-\nabla_{x}g(X_{s-}(\omega))\Delta X_{s}(\omega)\right)=M_{t}^{R}g(\omega)$
is an $(R,\mathbb{F})$-martingale as $R$ is a control rule in Definition 3.1. Moreover, we have, for $m\otimes R$-$\mathrm{a.s.}~(t,\omega)\in[0,T]\times\Omega$,
@@
$J(\alpha,\zeta)$ $=\mathbb{E}^{R}\left[\int_{0}^{T}f(t,X_{t},\mathcal{L}^{R}(X_{t}),\alpha_{t})\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}\right]$
$\leq\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}f(t,X_{t},R^{X}(t),u)q_{t}(\mathrm{d}u)\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}\right]=\Gamma(R).$
@@
We first show that $m\otimes R(A_{i}^{R})=T$ iff $R(\tilde{A}_{i}^{R})=1$. Here, recall that $A_{i}^{R}\subset[0,T]\times\Omega$ is defined by (5).
Necessity. For any $\omega\in\Omega$, define the $\omega$-section of $A_{i}^{R}$ that
@@
Then, it follows from Fubini’s Theorem that $T=m\otimes R(A_{i}^{R})=\int_{\Omega}m(A_{i}^{R}(\omega))R(\mathrm{d}\omega)$. This implies that there exists some $R$-null set $N\in\mathcal{F}$ such that $m(A_{i}^{R}(\omega))=T$ for all $\N\omega\in\Omega\backslash$ N. For such $\omega$ and $0\leq t_{1}<t_{2}\leq T$, it holds that
@@
Sufficiency. For $\omega\in\tilde{A}_{i}^{R}$, we have $m(A_{i}^{R}(\omega))=T$, and Fubini’s Theorem results in $m\otimes R(A_{i}^{R})=\int_{\Omega}m(A_{i}^{R}(\omega))R(\mathrm{d}\omega)=T$.
@@
For the last inequality, we used the fact that $\tilde{A}_{i}^{R}$ is closed and $R_{\ell}$ converges weakly to $R$ as $\ell\to\infty$. This yields that $R(\tilde{A}_{i}^{R})=1$, and hence $m\otimes R(\tilde{A}_{i}^{R})=T$, which completes the proof. ∎

### v3

@@ 5 Uniqueness and Stability of Solution to Constrained FBSDE
Hence, combining the above two inequalities leads to $\bar{X}=0,\bar{Y}=0$ and $\bar{Z}=0$. It then follows that $\bar{\alpha}=0$ and $\int_{0}^{\cdot}G(s)\mathrm{d}\bar{\zeta}_{s}=0$. Moreover, we have $\int_{0}^{t}A(s)\bar{\eta}_{s}\mathrm{d}s+\int_{0}^{t}B(s)\mathbb{E}\left[\bar{\eta}_{s}\right]\mathrm{d}s=0$ for all $t\in[0,T]$. Taking expectations on both sides of the above result, together with (C1) in Assumption 3, we conclude that $\mathbb{E}\left[\bar{\eta}_{t}\right]=0$, for $m$-a.e. $t\in[0,T]$, and hence $\int_{0}^{t}A(s)\bar{\eta}_{s}\mathrm{d}s=0$, for $m$-a.e. $t\in[0,T]$. This indicates that $\bar{\eta}=0$ in $\mathbb{L}_{\mathbb{F}}^{2,1}$ by utilizing Assumption 3-(C1). ∎
@@
    The constrained function $\phi$ is independent of the control variable and admits the linear form that $\phi(t,x,\mu)=A(t)x+B(t)\int_{\mathbb{R}^{n}}x\mu(\mathrm{d}x)+C(t)$ for $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$. Here, the coefficients $(A,B)\in C([0,T],\mathbb{R}^{n}\times\mathbb{R}^{n})$ and $C\in C([0,T];\mathbb{R})$ satisfying $|A(t)|\geq c_{0}>0$ and $|A(t)+B(t)|\geq c_{0}>0$ for some $c_{0}>0$ independent of $t\in[0,T]$.
  • (C5’)
    Set $\Psi(x,\mu):=\nabla_{x}g(x,\mu)+\int_{\mathbb{R}^{n}}\partial_{\mu}g(y,\mu)(x)\mu(\mathrm{d}y)$ for $(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$. The function $\Psi$ is Lipschitz continuous and satisfies the coercivity condition that, for some $\alpha\geq 0$ and $(x^{1},\mu^{1},X^{1}),(x^{2},\mu^{2},X^{2})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times L^{2}((\Omega,\mathcal{F}_{T},\mathbb{P}),\mathbb{R}^{n})$,
@@
Let Assumption 2, (C1’), (C3)-(C4) of Assumption 3 hold and $\xi_{1},\xi_{2}\in L^{2}((\Omega,\mathcal{F}_{0},\mathbb{P});\mathbb{R}^{n})$. Moreover, suppose that (C5’) holds for either $\Psi^{1}$ or $\Psi^{2}$. Denote by $(\widehat{X}^{i},\widehat{\alpha}_{t}^{i},\widehat{\zeta}^{i}_{t},Y^{i},Z^{i},\eta^{i})$ the unique solutions to the FBSDE (5)-(31) with respective initial data $\xi_{i}$. Set $\bar{X}_{t}:=X_{t}^{1}-X_{t}^{2}$, $\bar{\zeta}_{t}:=\zeta_{t}^{1}-\zeta_{t}^{2}$, $\bar{Y}_{t}:=Y_{t}^{1}-Y_{t}^{2}$, $\bar{Z}_{t}:=Z_{t}^{1}-Z_{t}^{2}$ and $\bar{\eta}_{t}:=\eta_{t}^{1}-\eta_{t}^{2}$ for $t\in[0,T]$. Then, there exists a constant $C>0$ depending on $(T,M,\alpha,\beta)$ only such that
@@
$\qquad\qquad\leq C\left\{\mathbb{E}\left[|\xi_{1}-\xi_{2}|^{2}\right]+\sup_{(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})}\left|\Psi^{1}(x,\mu)-\Psi^{2}(x,\mu)\right|^{2}\right\}$ (35)
@@
Recall $\Delta b(t)$, $\Delta\sigma(t)$ and $\Delta F(t)$ defined in the proof of Theorem 5.4. Then, one has
$\int_{0}^{t}G(s)\mathrm{d}\bar{\zeta}_{s}=\bar{X}_{t}-(\xi_{1}-\xi_{2})-\int_{0}^{t}\Delta b(s)\mathrm{d}s-\int_{0}^{t}\Delta\sigma(s){\mathrm{d}W_{s}}.$ (36)
@@
$\left|\bar{Y}_{t}\right|^{2}+\int_{t}^{T}|\bar{Z}_{s}|^{2}\mathrm{d}s$ $={|\bar{Y}_{T}|^{2}}+2\int_{t}^{T}\bar{Y}_{s}^{\top}\Delta F(s)\mathrm{d}s-2\int_{t}^{T}\bar{Y}_{s}\bar{Z}_{s}\mathrm{d}W_{s}-2\int_{t}^{T}\bar{\eta}_{s}A(s)^{\top}\bar{Y}_{s}\mathrm{d}s$
@@
Using Assumption 3-(C4), the condition (C1’) (C5’), the BDG’s inequality and the Gronwall’s inequality, we can similarly derive, for some constant $C>0$ depending on $(M,\epsilon,T)$ only,
@@
Applying integration by parts to $\bar{X}_{t}\bar{Y}_{t}$ from $t=0$ to $T$, we have
@@
Taking expectations on both sides of the above equality and utilizing (C5’) (we assume (C5’) holds for $\Psi^{1}$ without loss of generality), we deduce that
$\alpha\mathbb{E}\left[|\bar{X}_{T}|^{2}\right]+\beta\mathbb{E}\left[\int_{0}^{T}(|\bar{X}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2})\mathrm{d}t\right]$
$\quad\leq\mathbb{E}\left[\bar{Y}_{0}^{\top}\left(\xi^{1}-\xi^{2}\right)\right]+\sup_{(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})}\left|\Psi^{1}(x,\mu)-\Psi^{2}(x,\mu)\right|^{2}$
$\quad\leq\frac{2}{\epsilon}\mathbb{E}\left[\left|\bar{Y}_{0}\right|^{2}\right]+\frac{1}{2\epsilon}\mathbb{E}\left[\left|\xi^{1}-\xi^{2}\right|^{2}\right]+\sup_{(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})}\left|\Psi^{1}(x,\mu)-\Psi^{2}(x,\mu)\right|^{2}.$ (41)
@@
$\mathbb{E}\left[\left|\bar{X}_{T}\right|^{2}+\int_{0}^{T}\left(|\bar{X}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2}\right)\mathrm{d}t\right]$
$\quad\leq C\left(\mathbb{E}\left[\left|\xi^{1}-\xi^{2}\right|^{2}\right]+\sup_{(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})}\left|\Psi^{1}(x,\mu)-\Psi^{2}(x,\mu)\right|^{2}\right).$ (42)
Combining the Lipschitz continuity of $\hat{u}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto U\subset\mathbb{R}^{l}$ in Assumption 3-(C3), (36), (39) and the condition (C1’), we conclude the stability (35) from (5) above. ∎
@@
We recall that the existence of an optimal strict control is proved in Section 3 (see Theorem 3.10 and Corollary 3.11). The necessary conditions characterizing the optimal control are given in Theorem 4.2 using FJ condition, which in turn ensures the existence of solution to the associated FBSDE. Moreover, the uniqueness of the solution to the constrained FBSDE is established in Section 5. In sum, the combination of Theorem 3.10, the characterization results in Section 4 and the uniqueness result in Section 5 provide a complete theory for the existence and uniqueness of solution to the associated FBSDE and the characterization of the optimal control for the general constrained MFC problem.
Finally, we present an application example in monetary reserve management of interbank system similar to [12] and [5]. Consider the log-monetary reserve process of a representative bank within the interbank system under equilibrium governed by the following controlled McKean-Vlasov SDE:
$dX_{t}^{\zeta}=\alpha(\mathbb{E}[X_{t}^{\zeta}]-X_{t}^{\zeta})dt+\sigma dW_{t}+d\zeta_{t},\quad t\in[0,T].$ (43)
Here, the term $\alpha(\mathbb{E}[X_{t}^{\zeta}]-X_{t}^{\zeta})dt$ with $\alpha>0$ depicts that the bank’s reserves are constantly adjusting toward the average reserve level of the banking system. The term $\sigma dW_{t}$ with $\sigma>0$ describes the exogenous noise or randomness, representing market uncertainty affecting the bank’s liquidity, and the term $d\zeta_{t}$ stands for the singular control that captures the active intervention of the bank with the central bank. In particular, if $\zeta=(\zeta_{t})_{t\in[0,T]}$ is absolutely continuous w.r.t. $t$, then $\dot{\zeta}_{t}:=d\zeta_{t}/dt$ represents the rate of borrowing/lending to the central bank.
The goal of this MFC problem is to choose an optimal control $\zeta=(\zeta_{t})_{t\in[0,T]}$ that minimizes the cost functional that
$\inf_{\zeta\in\mathscr{U}_{\text{ad}}[0,T]}J(\zeta)=\inf_{\zeta\in\mathscr{U}_{\text{ad}}[0,T]}\mathbb{E}\left[\int_{0}^{T}\beta\left(\mathbb{E}[X_{t}^{\zeta}]-X_{t}^{\zeta}\right)^{2}dt+\int_{[0,T]}\gamma d\zeta_{t}\right].$ (44)
The objective functional includes two costs: (i) the deviation cost: the term $\beta(\mathbb{E}[X_{t}^{\zeta}]-X_{t}^{\zeta})^{2}$ penalizes the squared deviation of the individual bank’s reserves from the system average. A higher $\beta>0$ implies a stronger penalty for idiosyncratic risk or deviation from market norms, pushing the bank toward the mean; (ii) the trading cost: the term $\int_{0}^{T}\gamma d\zeta_{t}$ accounts for the cost of executing trades with the central bank. The parameter $\gamma\in\mathbb{R}$ determines whether borrowing from or lending to the central bank incurs a cost or generates profit.
The system is subject to the following dynamical constraint, which imposes a bound on the deviation between the expected reserve and the actual reserve:
@@
$\left|\kappa\mathbb{E}[X_{t}^{\zeta}]-X_{t}^{\zeta}\right|\leq c,\quad dt\times d\mathbb{P}\mbox{-}a.s.,$ (45)
where $c>0$ serves as a predetermined threshold value and the term $\kappa\mathbb{E}[X_{t}^{\zeta}]$ represents a scaled average of log‑monetary reserve with scaling factor $\kappa\in(0,1)$. The form of constraint in (45) reflects a tolerance buffer: a bank is allowed to deviate from the system average, but only up to a proportion $(1-\kappa)$ of that average. In practice, $\kappa$ can be interpreted as a regulatory confidence factor or a liquidity-adjustment coefficient. A lower $\kappa$ permits larger individual deviations from the mean, offering more flexibility.
From the model described by (43), (44) and (45), for $(x,\mu)\in\mathbb{R}\times{\cal P}_{2}(\mathbb{R})$, it holds that
$b(x,\mu)=\alpha\left(\int_{\mathbb{R}}y\mu(dy)-x\right),~~\sigma(x,\mu)\equiv\sigma,~~G(t)\equiv 1,~~f(x,\mu)=\beta\left(\int_{\mathbb{R}}y\mu(dy)-x\right)^{2},$
while $c(t)\equiv\gamma$ and the constraint functions are $\phi^{1}(x,\mu)=\kappa\int_{\mathbb{R}}y\mu(dy)-x+c$ and $\phi^{2}(x,\mu)=x-\kappa\int_{\mathbb{R}}y\mu(dy)+c$ in terms of dynamical constraint (45). It is not difficult to verify that the above model coefficients satisfy Assumption1 and Assumption 2. As the constraint functions $\phi^{1}(x,\mu)$ and $\phi^{2}(x,\mu)$ are independent of regular controls and satisfy Assumption 3-(C1), the KKT condition holds by Lemma 5.3. The Hamiltonian in Theorem 4.2 becomes that, for $(x,\mu,y,z)\in\mathbb{R}\times{\cal P}_{2}(\mathbb{R})\times\mathbb{R}^{2}$,
$H(x,\mu,y,z)=\alpha y\left(\int_{\mathbb{R}}y\mu(dy)-x\right)+\sigma z+\beta\left(\int_{\mathbb{R}}y\mu(dy)-x\right)^{2}.$
Let $\widehat{\zeta}=(\widehat{\zeta}_{t})_{t\in[0,T]}$ be the optimal singular control and $\widehat{X}=(\widehat{X}_{t})_{t\in[0,T]}$ be the state dynamics in (43) under $\widehat{\zeta}$. The SMP established in Theorem 4.2 gives that, the optimal singular control $\widehat{\zeta}=(\widehat{\zeta}_{t})_{t\in[0,T]}$ of the above C-MFC problem can be characterized by
$\mathbb{E}\left[\int_{0}^{T}\boldsymbol{1}_{\{Y_{t}+\gamma>0\}}\mathrm{d}\widehat{\zeta}_{t}\right]=0,$ (46)
where the pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ satisfies $Y_{t}\geq-\gamma$ for all $t\in[0,T]$, $\mathbb{P}$-a.s., and moreover, the pair of processes solves the constrained BSDE:
$\mathrm{d}Y_{t}$ $=\alpha(Y_{t}-\mathbb{E}[Y_{t}])\mathrm{d}t-2\beta(\widehat{X}_{t}-\mathbb{E}[\widehat{X}_{t}])\mathrm{d}t+(\widehat{\eta}_{t}^{1}-\kappa\mathbb{E}[\widehat{\eta}_{t}^{1}])\mathrm{d}t-(\widehat{\eta}_{t}^{2}-\kappa\mathbb{E}[\widehat{\eta}_{t}^{2}])\mathrm{d}t+Z_{t}dW_{t},$
$Y_{T}$ $=0.$ (47)
For this mean-field linear BSDE, the well-posedness can be addressed by Theorem 5.4. Additionally, the non-negative adapted square-integrable processes $\widehat{\eta}^{i}=(\widehat{\eta}_{t}^{i})_{t\in[0,T]}$ for $i=1,2$ satisfy that
$\mathbb{E}\left[\int_{0}^{T}\left(\kappa\mathbb{E}[\widehat{X}_{t}]-\widehat{X}_{t}+c\right)\widehat{\eta}_{t}^{1}\mathrm{d}t\right]=\mathbb{E}\left[\int_{0}^{T}\left(\widehat{X}_{t}-\kappa\mathbb{E}[\widehat{X}_{t}]+c\right)\widehat{\eta}_{t}^{2}\mathrm{d}t\right]=0.$ (48)
Remark 5.7.
Our C-MFC problem may be explained as the approximation of large-scale finite $N$-player cooperative game. An interesting future research is to establish the connection between the C-MFC problem and the N-player cooperative game under dynamical constraints. However, the main difficulty lies in the complex dynamic constraints in the N-player controlled system. In fact, to establish the connection of the approximation, one possible way is to transform the constrained N-player cooperative game into a problem without constraints. Then, we may apply the Lagrange-multiplier approach in [7] to derive the SMP for the $N$-player cooperative game problem. We then will encounter a system of coupled FBSDEs, which is in general very challenging. Moreover, the approximation or the propagation of chaos will further require the joint convergence of the associated Lagrange multipliers and singular controls. One may also need to verify that the limiting model still satisfies the dynamic constraints, which introduces additional analytical difficulties. We leave this interesting but challenging problem for future study.
@@
Acknowledgements The authors sincerely thank three anonymous referees for their constructive comments and suggestions, which significantly improve the manuscript. L. Bo is supported by National Natural Science Foundation of China (No. 12471451), Natural Science Basic Research Program of Shaanxi (No. 2023-JC-JQ-05), Shaanxi Fundamental Science Research Project for Mathematics and Physics (No. 23JSZ010). X. Yu and J. Wang are supported by the Hong Kong RGC General Research Fund (GRF) under grant No. 15214125 and by the Hong Kong Polytechnic University research grant under No. P0045654.
@@
This, together with Lemma 3.3, indicates that $\mathbb{E}^{R}[\int_{0}^{T}|\alpha_{t}|^{p}\mathrm{d}t]<\infty$. On the other hand, using the construction of $\alpha$ again, we deduce that, for any $\psi\in C_{b}^{2}(\mathbb{R}^{n})$ and $(t,\omega)\in[0,T]\times\Omega$,
$\psi(X_{t}(\omega))-\int_{0}^{t}\mathbb{G}\psi(s,X_{s}(\omega),\mathcal{L}^{R}(X_{s}),\alpha_{s}(\omega))\mathrm{d}s-\int_{0}^{t}\nabla_{x}\psi(X_{s}(\omega))G(s)\mathrm{d}\zeta_{s}(\omega)$
$\qquad-\sum_{0<s\leq t}\left(\psi(X_{s}(\omega))-\psi(X_{s-}(\omega))-\nabla_{x}\psi(X_{s-}(\omega))\Delta X_{s}(\omega)\right)=M_{t}^{R}\psi(\omega)$
is an $(R,\mathbb{F})$-martingale as $R$ is a control rule in Definition 3.1. Moreover, we have, for $m\times R$-$\mathrm{a.s.}~(t,\omega)\in[0,T]\times\Omega$,
@@
$J(\alpha,\zeta)$ $=\mathbb{E}^{R}\left[\int_{0}^{T}f(t,X_{t},\mathcal{L}^{R}(X_{t}),\alpha_{t})\mathrm{d}t+{\int_{[0,T]}c(t)\mathrm{d}\zeta_{t}+g(X_{T},\mathcal{L}^{R}(X_{T}))}\right]$
$\leq\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}f(t,X_{t},R^{X}(t),u)q_{t}(\mathrm{d}u)\mathrm{d}t+{\int_{[0,T]}c(t)\mathrm{d}\zeta_{t}+g(X_{T},\mathcal{L}^{R}(X_{T}))}\right]=\Gamma(R).$
@@
We first show that $(m\times R)(A_{i}^{R})=T$ iff $R(\tilde{A}_{i}^{R})=1$. Here, recall that $A_{i}^{R}\subset[0,T]\times\Omega$ is defined by (7).
Necessity. For any $\omega\in\Omega$, we define the $\omega$-section of $A_{i}^{R}$ that
@@
Then, it follows from Fubini’s Theorem that $T=(m\times R)(A_{i}^{R})=\int_{\Omega}m(A_{i}^{R}(\omega))R(\mathrm{d}\omega)$. This implies that, there exists some $R$-null set $N\in\mathcal{F}$ such that $m(A_{i}^{R}(\omega))=T$ for all $\N\omega\in\Omega\backslash$ N. For such $\omega$ and $0\leq t_{1}<t_{2}\leq T$, it holds that
@@
For the last inequality, we used the fact that $\tilde{A}_{i}^{R}$ is closed and $R_{\ell}$ converges weakly to $R$ as $\ell\to\infty$. This yields that $R(\tilde{A}_{i}^{R})=1$, and hence $(m\times R)(\tilde{A}_{i}^{R})=T$, which completes the proof. ∎